OpenStudy (anonymous):

Simplify the following expression: 5 times the square root of 12b minus 3 times the square root of 3b.

OpenStudy (anonymous):

so... \[5\sqrt{12b} - 3\sqrt{3b}\]?

OpenStudy (anonymous):

can you break down any of the terms to make them more similar?

OpenStudy (anonymous):

I'm not sure if you solve them separately then minus the two. or try like terms with it. My teacher never explained it to us well.

OpenStudy (anonymous):

well if you break down \[\sqrt{12} = \sqrt{2 \times 2 \times 3} = 2\sqrt{3}\] does that help?

OpenStudy (anonymous):

Yes thank you a bunch! (:

OpenStudy (anonymous):

no problem :)